# AVID: Any-Length Video Inpainting with Diffusion Model

### 06 Dec 2023 v1

@@ Abstract
Recent advances in diffusion models have successfully enabled text-guided image inpainting. While it seems straightforward to extend such editing capability into video domain, there has been fewer works regarding text-guided video inpainting. Given a video, a masked region at its initial frame, and an editing prompt, it requires a model to do infilling at each frame following the editing guidance while keeping the out-of-mask region intact. There are three main challenges in text-guided video inpainting: ($i$) temporal consistency of the edited video, ($ii$) supporting different inpainting types at different structural fidelity level, and ($iii$) dealing with variable video length. To address these challenges, we introduce Any-Length Video Inpainting with Diffusion Model, dubbed as AVID. At its core, our model is equipped with effective motion modules and adjustable structure guidance, for fixed-length video inpainting. Building on top of that, we propose a novel Temporal MultiDiffusion sampling pipeline with an middle-frame attention guidance mechanism, facilitating the generation of videos with any desired duration. Our comprehensive experiments show our model can robustly deal with various inpainting types at different video duration range, with high quality [^1] More visualization results is made publicly available here..
Figure 1: Video inpainting. We introduce a video inpainting method versatile across a spectrum of video durations and tasks. Displayed frames are uniformly selected from videos of different lengths. The first row in the figure contains the source videos and the target regions, while the bottom row shows the results. Caption in the middle represents the language guidance and duration for each video.
@@ 1 Introduction
Past years have witnessed remarkable advancements in image inpainting [37, 66, 10], which infills a given masked region in an image. Owing to recent explosive evolution of text-to-image (T2I) generative models, the content infilling becomes even more flexible by conditioning on a given textual description [66, 45]. On the other hand, text-to-video (T2V) generation is also developing rapidly [26]. This leads to a compelling inquiry: can we also harness this prowess for text-guided video inpainting?
@@
Imagine given a video of “a car traversing a road” and one wants to replace this car with a MINI Cooper. Perhaps one of the easiest editing approach is user just simply clicking on the car in the first frame and writing down a short textual prompt “MINI Cooper”, then an advanced model congests this information and generates a spatially-seamless and temporally-consistent moving MINI Cooper in the scene while not changing any other video parts. We believe such interaction will become a fundamental and popular video editing approach, thus decide to address this task as its pioneering work.
@@
There are mainly three challenges in video inpainting. First, the synthesized content must exhibit temporal consistency. If an entire object is rendered, its identity should persist throughout the video. For example, if a color of the car is changed to green, it is imperative that the hue remains consistent from start to finish, the car should remain the same shade of green, rather than transitioning from, say, neon green to a darker variant. Second, there are various editing types in video inpainting. For example, as in Fig. 1, one common inpainting type is object swap (e.g. “replace a car with a MINI Cooper”, another type is re-texturing (e.g. “turn the leaf color from red to yellow”), and uncropping (e.g. “fill in the upper and lower region”) is also popular. Different editing type requires a different level of structural fidelity to the original video. For instance, in object swap, if the editing is to transform a masked person into a statue, the structure and motion of the original person indicated by the mask region should be maintained in the output video. As comparison, video uncropping entails filling in blank spaces to augment the view, with no guidance signal from the mask region. Third, an input video can be of variable length, and thus we expect a good model should robustly handle any video duration.
@@
In light of these challenges, we propose Any-Length Video Inpainting with Diffusion Models, dubbed as AVID, a unified framework tailored for video inpainting. Built upon a text-guided image inpainting framework, we integrate motion modules to ensure the temporal coherence within the edited region. A structure guidance module is also incorporated, adaptable to the varying structural fidelity demands of different video inpainting tasks. Our innovation also includes a zero-shot generation pipeline, reinforced by an middle-frame attention guidance mechanism, enabling the handling of videos across varying durations. With the proposed model, given a textual description for the desired modification, along with a mask on the initial frame of a video indicating the area to edit as in Fig. 1, we can adeptly modify the content of variable video length, aligning them with the desired narrative while conserving intricate details, maintaining temporal consistency, and leaving the region outside the mask unaltered.
@@
To encapsulate our contributions: (I) We integrate motion modules [14] into a text-to-image inpainting model and optimize it on video sequences, thereby ensuring temporal consistency. (II) We propose a structure guidance module tailored for different sub-tasks, so that users can control different degrees of output’s structural fidelity towards the input video depending on the task and editing need. (III) We incorporate a pioneering zero-shot generation technique, proficiently handling different video length without additional training. Concurrently, we introduce an middle-frame attention guidance methodology, ensuring temporal consistency even in elongated video sequences. (IV) Rigorous evaluation on various inpainting tasks, region areas, and video length demonstrates the robustness of our model.
@@ 2 Related Work
The success of diffusion models [23, 53, 55] have enabled advanced image generation [44, 48, 16, 41, 43]. These models leverage large-scale text-image datasets [49] to produce remarkable outcomes. Further, recent works utilized the pre-trained text-to-image models for image manipulation in response to natural language descriptions [20, 12, 46, 27, 35, 32, 36, 38, 15, 19, 72, 18, 29, 54, 4, 6, 31]. Among them, customized image generation has particularly benefited from controlling the generated content via additional structure modules, e.g., ControlNet [71], T2I-adapter [40], etc.
Such techniques are also influencing a traditional but popular image editing task – image inpainting. While generative adversarial networks [69, 70] have been mainly applied for this task, recent diffusion-based models [47, 1] are showing more impressive results. However, these models are limited in filling in the content using out-of-mask context only. A more flexible usage is adding text control [68, 67, 3, 2, 60, 66] that allow for text-buided image inpainting. Latent Blended Diffusion [3] proposed blending the generated and original image latents, Imagenator [60] and Diffusion-based Inpainting [45] fine-tune pre-trained text-to-image generation models with masked images as additional input, and SmartBrush [66] fine-tunes an additional mask prediction branch on object-centric datasets.
Extending the success of text-guided image inpainting to video domain presents unique challenges, especially in maintaining the temporal consistency in videos of arbitrary duration. The scarcity of high-quality, large-scale video datasets [26, 59, 62, 63, 52, 17, 25, 5, 24, 14, 8] further complicates this task. Recent efforts have explored utilizing pre-trained image models for video editing, e.g. employing DDIM inversion [53] for consistent latents [50, 64, 42, 7, 61, 13]. Nevertheless, most approaches are proposed without considering an explicit mask input. Relying on the textual editing prompts only could easily change the undesired regions. As comparison, VideoComposer takes the masked frames as input and addresses video inpainting as one of their editing tasks, yet its constraint of applying a uniform target region across all frames limits its flexibility and sacrifices the editing quality.
@@ 3 Methods
For a video of arbitrary duration, a mask region, and a editing prompt, our objective is to fill in the indicated region at each frame following the editing guidance, while keeping out-of-mask video portion unchanged. We introduce our method, AVID, as depicted in Fig. 2. Our approach is built on top of a diffusion-based text-guided image inpainting model [45], then adapt it to video inpainting by inflating the base model and learning the motion modules (Sec. 3.2). We further add a structure guidance module to meet different inpainting type and structural fidelity requirement (Sec. 3.3). Additionally, we propose a zero-shot inference pipeline for dealing with videos of varying lengths (Sec. 3.4).
@@ 3.2 Text-guided Video Inpainting
Given a video $v_{0}=\{x_{0}^{i}\}_{i=1}^{N}$ and a mask region at its first frame $m^{0}$, we seek to edit the target region at each frame, aligning its content with the given text prompt $y$. In order to predict a precise editing region at each frame, we first propagate $m^{0}$ to every following frame and obtain a mask sequence $m=\{m^{i}\}_{i=1}^{N}$. This is achieved by applying XMem  [9] to track the masked region through the whole video, which we found to be simple and precise for editing type like object swap and re-texturing. For video uncropping, we simply use the same to-fill region for all frames.
To pursue a good temporal consistency across video frames, we follow AnimateDiff [14] to inflate an image diffusion model by converting its 2D layers to pseudo-3D and adding additional motion modules to learn the temporal correlations between frames. As described in  Eq. 1 to obtain $v_{t}$, the video frames undergo a forward diffusion process individually. Then our objective for video inpainting becomes:
@@
with $\mathbf{c}=\left(v_{m},m,\tau_{\theta}(y)\right)$ and $v_{m}=\{x_{0}^{i}\odot(1-m^{i})\}_{i=1}^{N}$. We optimize only the motion modules to retain the generative capabilities of the pre-trained image model, as illustrated in  Fig. 2 (a).
@@ 3.3 Structure Guidance in Video Inpainting
In practice, as in  Fig. 1, video inpainting involves different editing types, with respect to different structural fidelity. For example, re-texturing requires the structure of the source video to be preserved, e.g., converting the material of the person’s coat to leather, as in  Fig. 4, while uncropping has no such requirement. This inspires us to introduce an additional structure guidance module to our model. Following the design of ControlNet [71], we fix the parameters of $\epsilon_{\theta}$ and proceed to train the structure-conditioned module, denoted as $\mathbf{s}_{\theta}$, as in  Fig. 2 (b). We employ a structure information extractor $\mathcal{S}_{\theta}$ to obtain structure condition on each frame $s=\{s^{i}\}_{i=1}^{N}$, where $s^{i}=\mathcal{S}_{\theta}\left(x_{0}^{i}\right)$. The structure guidance module outputs $\mathbf{c}_{s}=\mathbf{s}_{\theta}\left(v_{t},t,\mathbf{c},s\right)$, composed of $13$ feature maps $\{h_{i}\}_{i=1}^{13}$ at $4$ different resolutions. These feature maps are then integrated into the skip connections and the output of the middle block of $\epsilon_{\theta}$, contributing to the generation process. Therefore, the training objective for this phase is formulated as:
@@ 3.4 Zero-shot Inference for Long Videos
Although motion modules (implemented as temporal self-attention layers) can take videos with varying number of frames theoretically, it suffers from drastic quality degradation when generating a longer video of more frames than training [14]. We show such quality degradation in supplementary material. Inspired by MultiDiffusion [4] generating a high-resolution image seamlessly consisting of multiple patches, we extend such idea into temporal axis, making it work effectively to deal with a video of any length. We further propose a novel middle-frame attention guidance mechanism, to keep the identity unchanged through out the video.
@@
where $v_{t-1}^{i}[j]$ is the processed frame from $v_{t}^{i}$ corresponding to $v_{t}^{\prime}[k]$, and $\lVert\mathbb{S}_{k}\rVert$ is the cardinality of the set $\mathbb{S}_{k}$, indicating the number of times frame $k$ is processed across different clips.
@@
Figure 4: Editing on videos of different durations. We employ our method on various videos and edit them on different tasks. We show the wide range of edits our approach can be used with different region sizes and video durations. Above each video, we note the number of frames, $N^{\prime}$, and the video duration.
@@ 4 Experiments
Figure 5: Comparison of various methods. We compare our method against several approaches, including per-frame inpainting using text-to-image LDM inpainting (PF) [45], Text2Video-Zero (T2V0) [28], VideoComposer (VC). All methods are evaluated using their default hyper-parameters as specified in either their corresponding publications or source codes. Each video in our experiments consists of $16$ frames. Our proposed approach successfully edits the videos as intended while retaining the details outside the designated target region. Moreover, our method upholds the editing capabilities of the image in-painting model we utilized. Notably, our results demonstrate remarkable consistency, outperforming other methods in our comparison.
Implementation details. Our implementation is built upon a large-scale pre-trained inpainting Latent Diffusion Model (LDM) [45]. For training, we use the watermark-removed Shutterstock video dataset [51], with motion modules being trained using $16$ frames at a $512\times 512$ resolution using synthetic random mask. Subsequently, the parameters from the UNet encoder are transferred to the control module, which is then trained using the same dataset. We use Holistically-Nested Edge Detection (HED)  [65] within in the synthetic region as the structure guidance for the control module training. During the structure guidance module training, all parameters in the control module are optimized.
@@ 4.2 Comparisons
Table 1: Quantitative results. We compare our method against several approaches, including per-frame in-painting (PF) using Stable Diffusion In-painting [45], Text2Video-Zero (T2V0) [28], and VideoComposer (VC) on different video generative fill sub-tasks and evaluate generated results using different metrics, including background preservation (BP ×$\times 10^{-3}$, $\downarrow$ better), text-video alignment (TA, $\uparrow$ better), and temporal consistency (TC, $\uparrow$ better). $*{}^{*}$ indicates structure guidance is applied for VC and our approach.
Qualitative comparisons. We present a comprehensive evaluation of our method against other diffusion-based video inpainting techniques, notably the per-frame inpainting techniques using inpainting LDM [45] and VideoComposer. We apply our model on the same videos as in [58], which have provided results of VideoComposer. Recognizing that video inpainting can be construed as a text-to-video generation with set boundary conditions, we also evaluate the training-free Text2Video-Zero [28] model on top of inpainting LDM as another comparison. To highlight the robustness of our foundational model, we limit our evaluations to videos spanning $16$ frames, intentionally omitting the zero-shot long video inferences introduced in  Sec. 3.4 for fair comparison.
Fig. 5 (left) compares the performance on re-texturing. Note both our model and VideoComposer can apply the structural guidance. We let VideoComposer integrate more cues (including the sketch maps, depth maps, and motion vectors) for its better results, while our model only applys HED. Fig. 5 (right) proceeds object swap without structural guidance for all methods. Specifically, the comparison between VideoComposer (row 4) and our model (row 5) reveals extraneous color shift in the former, whereas the latter demonstrates impeccable background preservation. Furthermore, compared with the other two zero-shot methodologies, per-frame editing (row 2) and Text2Video-Zero (row 3) [28], our method exhibits far better temporal consistency. For example, our model can consistently keep the synthesized purple car shape in Fig. 5, while Text2Video-Zero [28] presents temporal morphing inconsistencies.
Figure 6: User study. In our user preference studies, we juxtaposed our method against per-frame in-painting techniques by evaluating prominent models such as Diffusion-based Image In-painting[45], Text2Video-Zero (T2V0)[28], and VideoComposer (VC), assessing their performances across various tasks.

### 20 Mar 2024 v2

@@ Abstract
Recent advances in diffusion models have successfully enabled text-guided image inpainting. While it seems straightforward to extend such editing capability into the video domain, there have been fewer works regarding text-guided video inpainting. Given a video, a masked region at its initial frame, and an editing prompt, it requires a model to do infilling at each frame following the editing guidance while keeping the out-of-mask region intact. There are three main challenges in text-guided video inpainting: ($i$) temporal consistency of the edited video, ($ii$) supporting different inpainting types at different structural fidelity levels, and ($iii$) dealing with variable video length. To address these challenges, we introduce Any-Length Video Inpainting with Diffusion Model, dubbed as AVID. At its core, our model is equipped with effective motion modules and adjustable structure guidance, for fixed-length video inpainting. Building on top of that, we propose a novel Temporal MultiDiffusion sampling pipeline with a middle-frame attention guidance mechanism, facilitating the generation of videos with any desired duration. Our comprehensive experiments show our model can robustly deal with various inpainting types at different video duration ranges, with high quality [^1] More visualization results are made publicly available here..
Figure 1: Video inpainting. We introduce a video inpainting method versatile across a spectrum of video durations and tasks. Displayed frames are uniformly selected from videos of different lengths. The first row in the figure contains the source videos and the target regions, while the bottom row shows the results. The caption in the middle represents the language guidance and duration for each video.
@@ 1 Introduction
Past years have witnessed remarkable advancements in image inpainting [39, 71, 10], which infills a given masked region in an image. Owing to the recent explosive evolution of text-to-image (T2I) generative models, the content infilling becomes even more flexible by conditioning on a given textual description [71, 48]. On the other hand, text-to-video (T2V) generation is also developing rapidly [26]. This leads to a compelling inquiry: can we also harness this prowess for text-guided video inpainting?
@@
Imagine being given a video of “a car traversing a road” and one wants to replace this car with a MINI Cooper. Perhaps one of the easiest editing approaches is the user just simply clicking on the car in the first frame and writing down a short textual prompt “MINI Cooper”, then an advanced model congests this information and generates a spatially-seamless and temporally consistent moving MINI Cooper in the scene while not changing any other video parts. We believe such interaction will become a fundamental and popular video editing approach, thus decide to address this task as its pioneering work.
@@
There are mainly three main challenges in video inpainting. First, the synthesized content must exhibit temporal consistency. If an entire object is rendered, its identity should persist throughout the video. For example, if the color of the car is changed to green, it is imperative that the hue remains consistent from start to finish, the car should remain the same shade of green, rather than transitioning from, say, neon green to a darker variant. Second, there are various editing types in video inpainting. For example, as in Fig. 1, one common inpainting type is object swap (e.g. “replace a car with a MINI Cooper”, another type is re-texturing (e.g. “turn the leaf color from red to yellow”), and uncropping (e.g. “fill in the upper and lower region”) is also popular. Different editing types require different levels of structural fidelity to the original video. For instance, in object swap, if the editing is to transform a masked person into a statue, the structure and motion of the original person indicated by the mask region should be maintained in the output video. In comparison, video uncropping entails filling in blank spaces to augment the view, with no guidance signal from the mask region. Third, an input video can be of variable length, and thus, we expect a good model should robustly handle any video duration.
@@
In light of these challenges, we propose Any-Length Video Inpainting with Diffusion Model, dubbed as AVID, a unified framework tailored for video inpainting. Built upon a text-guided image inpainting framework, we integrate motion modules to ensure temporal coherence within the edited region. A structure guidance module is also incorporated, adaptable to the varying structural fidelity demands of different video inpainting tasks. Our innovation also includes a zero-shot generation pipeline, reinforced by a middle-frame attention guidance mechanism, enabling the handling of videos across varying durations. With the proposed model, given a textual description for the desired modification, along with a mask on the initial frame of a video indicating the area to edit as in Fig. 1, we can adeptly modify the content of variable video length, aligning them with the desired narrative while conserving intricate details, maintaining temporal consistency, and leaving the region outside the mask unaltered.
@@
To encapsulate our contributions: (I) We integrate motion modules [14] into a text-to-image inpainting model and optimize it on video sequences, thereby ensuring temporal consistency. (II) We propose a structure guidance module tailored for different sub-tasks, so that users can control different degrees of output’s structural fidelity towards the input video depending on the task and editing need. (III) We incorporate a pioneering zero-shot generation technique, proficiently handling different video lengths without additional training. Concurrently, we introduce a middle-frame attention guidance methodology, ensuring temporal consistency even in elongated video sequences. (IV) Rigorous evaluation on various inpainting tasks, region areas, and video length demonstrates the robustness of our model.
@@ 2 Related Work
The success of diffusion models [23, 56, 58] have enabled advanced image generation [47, 51, 16, 43, 46]. These models leverage large-scale text-image datasets [52] to produce remarkable outcomes. Further, recent works utilized the pre-trained text-to-image models for image manipulation in response to natural language descriptions [20, 12, 49, 27, 37, 32, 38, 40, 15, 19, 34, 33, 79, 18, 57, 4, 59, 6, 31]. Among them, customized image generation has particularly benefited from controlling the generated content via additional structure modules, e.g., ControlNet [78], T2I-adapter [42], etc.
Such techniques are also influencing a traditional but popular image editing task – image inpainting. While generative adversarial networks [75, 76] have been mainly applied for this task, recent diffusion-based models [50, 1] are showing more impressive results. However, these models are limited in filling in the content using out-of-mask context only. A more flexible usage is adding text control [74, 73, 3, 2, 64, 71] that allow for text-guided image inpainting. Latent Blended Diffusion [3] proposed blending the generated and original image latents, Imagenator [64] and Diffusion-based Inpainting [48] fine-tune pre-trained text-to-image generation models with masked images as additional input, and SmartBrush [71] fine-tunes an additional mask prediction branch on object-centric datasets.
Extending the success of text-guided image inpainting to video domain presents unique challenges, especially in maintaining the temporal consistency in videos of arbitrary duration. The scarcity of high-quality, large-scale video datasets [26, 63, 67, 68, 55, 17, 25, 5, 24, 14, 8] further complicates this task. Recent efforts have explored utilizing pre-trained image models for video editing, e.g. employing DDIM inversion [56] for consistent latents [53, 69, 45, 7, 65, 13]. Nevertheless, most approaches are proposed without considering an explicit mask input. Relying on the textual editing prompts only could easily change the undesired regions. As comparison, VideoComposer [66] takes the masked frames as input and addresses video inpainting as one of their editing tasks, yet its constraint of applying a uniform target region across all frames limits its flexibility and sacrifices the editing quality.
@@ 3 Methods
For a video of arbitrary duration, a mask region, and an editing prompt, our objective is to fill in the indicated region at each frame following the editing guidance, while keeping the out-of-mask video portion unchanged. We introduce our method, AVID, as depicted in Fig. 2. Our approach is built on top of a diffusion-based text-guided image inpainting model [48], then adapt it to video inpainting by inflating the base model and learning the motion modules (Sec. 3.2). We further add a structure guidance module to meet different inpainting types and structural fidelity requirements (Sec. 3.3). Additionally, we propose a zero-shot inference pipeline to deal with videos of varying lengths (Sec. 3.4).
@@ 3.2 Text-guided Video Inpainting
Given a video $v_{0}=\{x_{0}^{i}\}_{i=1}^{N}$ and a mask region at its first frame $m^{0}$, we seek to edit the target region at each frame, aligning its content with the given text prompt $y$. In order to predict a precise editing region at each frame, we first propagate $m^{0}$ to every following frame and obtain a mask sequence $m=\{m^{i}\}_{i=1}^{N}$. This is achieved by applying XMem  [9] to track the masked region through the whole video, which we found simple and precise for editing types like object swap and re-texturing. For uncropping, we simply use the same to-fill region for all frames.
To pursue good temporal consistency across video frames, we follow AnimateDiff [14] to inflate an image diffusion model by converting its 2D layers to pseudo-3D and adding additional motion modules to learn the temporal correlations between frames. As described in  Eq. 1 to obtain $v_{t}=\{x_{t}^{i}\}_{i=1}^{N}$, the video frames undergo a forward diffusion process individually. Then our objective for video inpainting becomes:
@@
with $\mathbf{c}=\left(v_{m},m,\tau_{\theta}(y)\right)$ and $v_{m}=\{x_{0}^{i}\odot(1-m^{i})\}_{i=1}^{N}$. We optimize only the motion modules to retain the generative capabilities of the pre-trained model, as in  Fig. 2 (a).
@@ 3.3 Structure Guidance in Video Inpainting
In practice, as in  Fig. 1, video inpainting involves different editing types, with respect to different structural fidelity. For example, re-texturing requires the structure of the source video to be preserved, e.g., converting the material of the person’s coat to leather, as in  Fig. 4, while uncropping has no such requirement. This inspires us to introduce an additional structure guidance module to our model. Following the design of ControlNet [78], we fix the parameters of $\epsilon_{\theta}$ and proceed to train the structure-conditioned module, denoted as $\mathbf{s}_{\theta}$, as in  Fig. 2 (b). We employ a structure information extractor $\mathcal{S}$ to obtain structure condition on each frame $s=\{s^{i}\}_{i=1}^{N}$, where $s^{i}=\mathcal{S}\left(x_{0}^{i}\right)$. The structure guidance module outputs $\mathbf{c}_{s}=\mathbf{s}_{\theta}\left(v_{t},t,\mathbf{c},s\right)$, composed of $13$ feature maps $\{h_{i}\}_{i=1}^{13}$ at $4$ different resolutions. These feature maps are then integrated into the skip connections and the output of the middle block of $\epsilon_{\theta}$, contributing to the generation process. Therefore, the training objective for this phase is formulated as:
@@ 3.4 Zero-shot Inference for Long Videos
Although motion modules (implemented as temporal self-attention layers) can take videos with varying numbers of frames theoretically, it suffers from drastic quality degradation when generating a longer video of more frames than training [14]. We show such quality degradation in supplementary material. Inspired by MultiDiffusion [4] generating a high-resolution image seamlessly consisting of multiple patches, we extend such idea into the temporal axis, making it work effectively to deal with a video of any length. We further propose a novel middle-frame attention guidance mechanism, to keep the identity unchanged throughout the video.
@@
where $v_{t-1}^{i}[j]$ is the processed frame from $v_{t}^{i}$ corresponding to $v_{t}^{\prime}[k]$, and $\lVert\mathbb{S}_{k}\rVert$ is the cardinality of the set $\mathbb{S}_{k}$, indicating the number of times frame $k$ is processed across all clips.
@@
Figure 4: Editing on videos of different durations. We employ our method on various videos and edit them for different tasks. We show the wide range of edits our approach can be used with different region sizes and video durations. Above each video, we note the number of frames, $N^{\prime}$, and the video duration.
@@ 4 Experiments
Figure 5: Comparison of various methods. We compare our method against several approaches, including per-frame inpainting using text-to-image LDM inpainting (PF) [48], Text2Video-Zero (T2V0) [28], VideoComposer (VC) [66]. All methods are evaluated using their default hyper-parameters as specified in either their corresponding publications or source codes. Each video in our experiments consists of $16$ frames. Our proposed approach successfully edits the videos as intended while retaining the details outside the designated target region. Moreover, our method upholds the editing capabilities of the image in-painting model we utilized. Notably, our results demonstrate remarkable consistency, outperforming other methods in our comparison.
Implementation details. Our implementation is built upon a large-scale pre-trained inpainting Latent Diffusion Model (LDM) [48]. For training, we use the watermark-removed Shutterstock video dataset [54], with motion modules being trained using $16$ frames at a $512\times 512$ resolution using synthetic random mask. Subsequently, the parameters from the UNet encoder are transferred to the control module, which is then trained using the same dataset. We use Holistically-Nested Edge Detection (HED)  [70] within the synthetic region as the structure guidance for the control module training. During the structure guidance module training, all parameters in the control module are optimized.
@@ 4.2 Comparisons
Table 1: Quantitative results. We compare our method against several approaches, including per-frame in-painting (PF) using Stable Diffusion In-painting [48], Text2Video-Zero (T2V0) [28], and VideoComposer (VC) [66] on different video inpainting sub-tasks and evaluate generated results using different metrics, including background preservation (BP ×$\times 10^{-3}$, $\downarrow$ better), text-video alignment (TA, $\uparrow$ better), and temporal consistency (TC, $\uparrow$ better). $*{}^{*}$ indicates structure guidance is applied for VC and our approach.
Qualitative comparisons. We present a comprehensive evaluation of our method against other diffusion-based video inpainting techniques, notably the per-frame inpainting techniques using inpainting LDM [48] and VideoComposer [66]. We apply our model on the same videos as in [62], which have provided results of VideoComposer [66]. Recognizing that video inpainting can be construed as a text-to-video generation with set boundary conditions, we also evaluate the training-free Text2Video-Zero [28] model on top of inpainting LDM as another comparison. To highlight the robustness of our foundational model, we limit our evaluations to videos spanning $16$ frames, intentionally omitting the zero-shot long video inferences introduced in  Sec. 3.4 for fair comparison.
Fig. 5 (left) compares the performance on re-texturing. Note both our model and VideoComposer [66] can apply the structural guidance. We let VideoComposer [66] integrate more cues (including the sketch maps, depth maps, and motion vectors) for better results, while our model only applies HED. Fig. 5 (right) proceeds object swap without structural guidance for all methods. Specifically, the comparison between VideoComposer [66] (row 4) and our model (row 5) reveals an extraneous color shift in the former, whereas the latter demonstrates impeccable background preservation. Furthermore, compared with the other two zero-shot methodologies, per-frame editing (row 2) and Text2Video-Zero (row 3) [28], our method exhibits far better temporal consistency. For example, our model can consistently keep the synthesized purple car shape in Fig. 5, while Text2Video-Zero [28] presents temporal inconsistencies.
Figure 6: User study. In our user preference studies, we juxtaposed our method against per-frame in-painting techniques by evaluating prominent models such as Diffusion-based Image In-painting[48], Text2Video-Zero (T2V0)[28], and VideoComposer (VC)[66], assessing their performances across various tasks.